ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 68 Grayson, Daniel, University of Illinois, USA Macaulay 2 Macaulay 2 is a software system designed to support research in algebraic geometry and commutative algebra. It provides fast algorithms for computing Groebner bases and resolutions of modules and an object oriented language for the user that allows high-level mathematical concepts to be implemented easily. We will demonstrate of the software, showing some of its main features. At the web site http://www.math.uiuc.edu/Macaulay2 you may preview the documentation or download the software.